OpenStudy (anonymous):

indefinite integral of 1/(3+25x^2)dx

OpenStudy (lgbasallote):

go @Mimi_x3 !!! she's awesome ;)

OpenStudy (mimi_x3):

\[\large \int\limits\frac{1}{3+25x^{2}} dx\] You can either use trig-sub or table of standard integral? Which one do you want to use?

OpenStudy (mimi_x3):

@lgbasallote is awesomeerrr!

OpenStudy (anonymous):

could you possibly do the beginnings of both? this is final review so i just need a refresher

OpenStudy (lgbasallote):

partial fractions is fun....lol

OpenStudy (mimi_x3):

well, if you want to use the table of standard integrals you can use: \[\int\limits\frac{1}{25(\frac{3}{25}+x^{2})} => \frac{1}{25} \int\limits\frac{1}{\left(\frac{\sqrt{3}}{5}\right)^{2}+x^{2}} \]

OpenStudy (mimi_x3):

And trig-sub is: \(x=\frac{\sqrt{3}}{5} tan\theta\)

OpenStudy (lgbasallote):

or use trig-sub from the start..makes no difference

OpenStudy (mimi_x3):

No, trig sub or table of standard integrals. but trig-sub is easier.